# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

### v1

@@ 2 Formulation
$\widetilde{D}=\{W_{2}\in\mathbb{R}^{n\times m}\colon W_{2}={\rm blocdiag}\{W_{2,1},\dots,W_{2,n}\}\}$
with $W_{2,j}\in\mathbb{R}^{1\times D_{j}}$ for all $j=1,\dots,n$. It is demonstrate that, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions (see Definition 4, Definition 5 and Definition 8 of [52]), then $D\cap\widetilde{D}$ (a subset of $D$) is nonempty (see Proposition 1-3 of [52]). Hence, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions, problem (8) is feasible. Combining with the fact that $B_{1}B_{1}^{\top}\succ 0$ and $W_{3}$ is a free variable to be chosen such that $W\succ 0$ (utilizing Schur complement lemma), problem (8) is strictly feasible, i.e., Assumption 3 holds. It is important to note that the feasible set of (8) is larger than the ones in [32, 49, 18, 52], since there is no structural constraints on $W_{2}$ in this paper.
@@
In fact, let
$D_{k}=\begin{bmatrix}0\\ I_{k}\end{bmatrix}\in\mathbb{R}^{n\times k},$
then the $\frac{n(n-1)}{2}$ equality constraints
$V_{j1}WV_{j2}=0,j=1,\dots,\frac{n(n-1)}{2}$ (9)
can be converted into $n-1$ equality constraint:
$e_{k}V_{2}WV_{2}^{\top}D_{n-k}=0,k=1,\dots,n-1$
with $e_{k}=(0,\dots,1,\dots,0)$. For the convenience and simplicity of formulas’ deducing, we still use $\frac{n(n-1)}{2}$ equality constraints (i.e., (9)) to describe the diagonal property of matrix $W_{1}$.
In Section 3, we will introduce a two-timescale algorithm framework based on differential equation solver approach to address optimization problem (8), which encompasses two distinct subproblems with one possessing closed form solution. The other subproblem is investigated in Section 4, and Section 5 provides the convergence analysis of the proposed algorithm and studies an acceleration optimization problem. In Section 6, we will study an $\ell_{0}$ sparsity promoting optimization problem that differs from (8), and propose some theorems to characterize its theoretical features.
@@ 3 Two-Timescale Optimization Algorithm Framework
Review the problem (8)
$\min_{W,P}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle R,W\rangle+\gamma g(P)$
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{G}(W)\in\mathcal{K},$
$V_{1}WV_{2}^{\top}-P=0.$
Let $\widetilde{W}={\rm vec}(W),\widetilde{P}={\rm vec}(P)$, then (8) is equivalent to the following optimization problem
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\gamma g(P)$ (10)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,\dots,M,$
@@
where $\widetilde{\Psi}_{i}={\rm vec}(\Psi_{i})$ and
$\Gamma_{+}^{p}=\left\{{\rm vec}(W)\colon W\in\mathbb{S}_{+}^{p}\right\}.$
If denote $\mathcal{A}\in\mathbb{R}^{(M+mn)\times p^{2}},\mathcal{B}\in\mathbb{R}^{(M+mn)\times mn}$ with
$\mathcal{A}=\begin{pmatrix}V_{12}^{\top}\otimes V_{11}\\ \vdots\\ V_{M2}^{\top}\otimes V_{M1}\\ V_{2}\otimes V_{1}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}=\begin{pmatrix}0\\ \vdots\\ 0\\ -I_{mn}\end{pmatrix},$
then optimization problem (10) is summarized in the following form
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $f(\widetilde{W})+h(\widetilde{P})$ (11)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0,$
@@
$f(\widetilde{W})$ $=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}),$
$h(\widetilde{P})$ $=\gamma g(P).$
To solve optimization problem (11), we briefly review the basic setup of primal-dual splitting method based on differential equation approach; see [31] for a comprehensive review. The following separable convex optimization problem is considered in [31]:
$\min_{x\in\mathcal{X},y\in\mathcal{Y}}F(x,y)=f(x)+g(y)\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm s.t.}\leavevmode\nobreak\ Ax+By=b,$ (12)
where $\mathcal{X}\subset\mathbb{R}^{m}$ and $\mathcal{Y}\subset\mathbb{R}^{n}$ are two closed convex sets, and $f$ and $g$ are two properly closed convex functions (not necessary differential). Assume that $g\in\mathcal{S}_{\mu_{g}}^{0}(\mathcal{Y})$ with $\mu_{g}>0$ and $f=f_{1}+f_{2}$ with $f_{2}\in\mathcal{S}_{0}^{0}(\mathcal{X})$, $f_{1}\in\mathcal{S}_{\mu_{f},L_{f}}^{1,1}(\mathcal{X})$ and $0\leq\mu_{f}\leq L_{f}<\infty$. Then, the following scheme is proposed to solve the above equality constrained optimization problem
[…]
@@
$\tilde{y}_{k}$ $=y_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-y_{k}),$ (15)
$\tilde{\eta}_{f,k}$ $=\gamma_{k}+\mu_{f}\alpha_{k},$ (16)
$\tilde{v}_{k}$ $=\frac{1}{\tilde{\eta}_{f,k}}(\gamma_{k}v_{k}+\mu_{f}\alpha_{k}u_{k}).$ (17)
To emphasis, $\alpha_{k}>0$ is the step size and the parameter system is discretized implicitly by
$\frac{\theta_{k+1}-\theta_{k}}{\alpha_{k}}=-\theta_{k+1},\quad\frac{\gamma_{k+1}-\gamma_{k}}{\alpha_{k}}=\mu_{f}-\gamma_{k+1},\quad\frac{\beta_{k+1}-\beta_{k}}{\alpha_{k}}=\mu_{g}-\beta_{k+1}$ (18)
@@
is the continuous-time version of algorithm (13), where $\mathcal{L}$ is defined in (39) below (see Section 5). It can be observed that differential inclusion (19) encompasses two Nesterov-type differential inclusions (studied in [47]) pertaining to primal variables $x$ and $y$, as well as a gradient flow about the dual variable $\lambda$. Besides, an alternative presentation of (19) reads as
@@
Therefore, (13) represents the semi-implicit Euler discretization of differential inclusion (20), which justifies its classification as a differential equation solver approach.
@@
We first consider the iteration of variable $\widetilde{P}\in\mathbb{R}^{mn\times 1}$. By (13),
$\widetilde{P}_{k+1}$ $=\mathbf{prox}_{\tau_{k}h}^{\mathbb{R}^{mn\times 1}}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right)$ (21)
$=\left(\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\sum_{i,j}\left(\gamma w_{ij}|P_{ij}|+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right)\right\}\right)_{mn\times 1},$ (22)
where $\rho_{k}=\frac{1}{\tau_{k}}$ and
$\Omega^{k}={\operatorname*{matrix}}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right).$ (23)
Letting $\theta(\widetilde{P})=\sum_{i,j}\left(\gamma w_{ij}|P_{ij}|+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right)$, it is obvious that
$\partial_{P_{ij}}\theta(\widetilde{P})=\left\{\begin{aligned} &\gamma w_{ij}[-1,1]+\rho_{k}\left(P_{ij}-\Omega^{k}_{ij}\right),\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}=0,\\ &\gamma w_{ij}{\operatorname*{sign}}(P_{ij})+\rho_{k}\left(P_{ij}-\Omega^{k}_{ij}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm otherwise}.\end{aligned}\right.$
Denoting $a_{k}=\frac{\gamma}{\rho_{k}}w_{ij}$, we have
$P_{ij}^{*}=\left\{\begin{aligned} &\left(1-\frac{a_{k}}{|\Omega_{ij}^{k}|}\right)\Omega_{ij}^{k},\leavevmode\nobreak\ \leavevmode\nobreak\ &|\Omega_{ij}^{k}|>a_{k},\\ &0,\leavevmode\nobreak\ \leavevmode\nobreak\ &|\Omega_{ij}^{k}|\leq a_{k}.\end{aligned}\right.$ (24)
Hence, $\widetilde{P}$ has naturally sparsity structure.
@@ 4 Solving Inner Layer Optimization Problem
Algorithm (13) consists of two subproblems: updating $v_{k+1}$ and $y_{k+1}$. The solution to the latter can be obtained explicitly using equation (24), while the solution to the former will be discussed in this section. We first deduce the Lagrange dual problem of the above subproblem due to the appealing structure of dual problem. Then, the BSUM framework is utilized to address the dual problem. Observe that the process of BSUM entails the sequential resolution of various optimization issues. However, each problem has a predefined structure, necessitating the need for individual design. Via the special and ingenious construction, each issue possesses a closed form solution and BSUM exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate.
The update of $v_{k+1}$ is given by
$v_{k+1}=\underset{v\in\mathcal{X}}{\operatorname*{argmin}}\left\{f_{2}(v)+\langle d_{k},v\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|Av+Bw_{k}-b\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|v-\widetilde{v}_{k}\right\|^{2}\right\},$ (25)
and when employing the symbol system of this paper, optimization problem (25) is transformed to
$\min_{{\rm vec}(W)}\leavevmode\nobreak$ $\langle d_{k},{\rm vec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}{\rm vec}(W)+\mathcal{B}w_{k}\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|{\rm vec}(W)-\widetilde{v}_{k}\right\|^{2}$ (26)
${\rm s.t.}\leavevmode\nobreak$ ${\rm vec}(W)\in\Gamma_{+}^{p},$
${\rm vec}(\Psi_{i})\in\Gamma_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
where
${\rm vec}(\Psi_{i})$ $={\rm vec}(-VF_{i}WV^{\top}-VWF_{i}^{\top}V^{\top}-VQV^{\top})$
$=-\left((V\otimes VF_{i})+(VF_{i}\otimes V)\right){\rm vec}(W)-(V\otimes V){\rm vec}(Q).$
By Assumption 3, for all $k$, problem (26) satisfies Slater’s condition and strong duality holds. Since $W,\Psi_{i}$ are symmetric matrices, there exist matrices $T_{1}\in\mathbb{R}^{\frac{p(p+1)}{2}\times p^{2}}$ and $T_{2}\in\mathbb{R}^{\frac{n(n+1)}{2}\times n^{2}}$ such that
${\rm svec}(W)$ $=T_{1}{\rm vec}(W)\leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ T_{1}T_{1}^{\top}=I_{\frac{p(p+1)}{2}},$
${\rm svec}(\Psi_{i})$ $=T_{2}{\rm vec}(\Psi_{i})\leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ T_{2}T_{2}^{\top}=I_{\frac{n(n+1)}{2}},$
where svec is the vectorization operator on $\mathbb{S}^{n}$ defined by
@@
We partition the identity matrix $I_{\frac{n(n+1)}{2}}$ as follows
$I_{\frac{n(n+1)}{2}}=(u_{11},u_{21},\dots,u_{n1},u_{22},\dots,u_{n2},\dots,u_{nn}).$
@@
with $E_{ij}=e_{i}e_{j}^{\top}$. By Lemma 3.2 of [33], $L^{\top}=L^{\dagger}$, and naturally $LL^{\top}=I_{\frac{n(n+1)}{2}}$.
Remark 7.
Inversely, there exist matrix $D_{1}\in\mathbb{R}^{p^{2}\times\frac{p(p+1)}{2}}$ and $D_{2}\in\mathbb{R}^{n^{2}\times\frac{n(n+1)}{2}}$ such that
${\rm vec}(W)=D_{1}{\rm svec}(W),\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm vec}(\Psi_{i})=D_{2}{\rm svec}(\Psi_{i}).$
@@
Then, by utilizing svec operator and denoting ($\widetilde{\Gamma}_{+}^{n}$ is defined analogously)
@@
$\min_{{\rm svec}(W)}\leavevmode\nobreak$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$ (27)
${\rm s.t.}\leavevmode\nobreak$ ${\rm svec}(W)\in\widetilde{\Gamma}_{+}^{p}$
${\rm svec}(\Psi_{i})\in\widetilde{\Gamma}_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
where ${\rm svec}(\Psi_{i})=T_{2}{\rm vec}(\Psi_{i}).$ The prime problem of (27) presents a level of difficulty that warrants exploration of its dual problem. The Lagrangian of (27) is defined as
$\mathcal{L}({\rm svec}(W);X)=$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$ (28)
$-\left\langle X_{0},{\rm svec}(W)\right\rangle+\left\langle(X_{1}+\dots+X_{M}),T_{2}(V\otimes V){\rm vec}(Q)\right\rangle$
@@
$\tilde{d}_{k}=D_{1}^{\top}d_{k}-X_{0}+D_{1}^{\top}\sum_{i=1}^{M}\left[(V\otimes VF_{i})^{\top}+(VF_{i}\otimes V)^{\top}\right]T_{2}^{\top}X_{i}\doteq\mathcal{L}_{k}(X_{0},X_{1},\dots,X_{M}),$
$\sigma_{1}=\frac{\alpha_{k}}{2\theta_{k}},\leavevmode\nobreak\ \sigma_{2}=\frac{\tilde{\eta}_{f,k}}{2\alpha_{k}},\leavevmode\nobreak\ \tilde{b}_{k}=\mathcal{B}w_{k},$

### v3

@@ 2 Formulation
$\widetilde{D}=\left\{W_{2}\in\mathbb{R}^{n\times m}\colon W_{2}=\mathrm{blocdiag}\{W_{2,1},\dots,W_{2,p}\}\right\}$
with $W_{2,j}\in\mathbb{R}^{n_{j}\times m_{j}}$ for all $j=1,\dots,p$ and $\sum_{j=1}^{p}m_{j}=m$, $\sum_{j=1}^{p}n_{j}=n$. It is demonstrated that, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions (see Definition 4, Definition 5 and Definition 8 of [14]), then $D\cap\widetilde{D}$ (a subset of $D$) is nonempty (see Proposition 1-3 of [14]). Hence, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions, problem (15) is feasible. Combining with the fact that $B_{1}B_{1}^{\top}\succ 0$ and $W_{3}$ is a free variable to be chosen such that $W\succ 0$ (utilizing Schur complement lemma), problem (15) is strictly feasible, i.e., Assumption 3 holds. It is important to note that the feasible set of (15) is larger than the ones in [40, 17, 47, 14], since there is no structural constraints on $W_{2}$ in this paper.
@@
By modifying optimization problem (15) as
$\min_{W,P}{}{}$ $\langle R,W\rangle$ (16)
${\rm s.t.}{}{}$ $\mathcal{G}(W)\in\mathcal{K},$
$[V_{1}WV_{2}^{\top}]_{ij}=0,~{}(i,j)\in\mathcal{U}\subseteq[m]\times[n],$
with $[\cdot]_{ij}$ the $(i,j)$-th element of a matrix and $\mathcal{U}$ a given subset, the methods proposed in this paper can be used to decentralized LQ problems with fixed communication topology, which have been widely studied [15, 16, 41, 42]. However, our paper is of independent interests for problems with fixed communication topology. Specifically, this paper provides an optimal-guarantee w.r.t. (16) rather than heuristic method [15] or stationary-point-guarantee [16]. Additionally, the proposed problem (16) is finite-dimensional, in comparison to the infinite-dimensional optimization [41, 42]. Moreover, if we only hope that agent $i_{k}$ cannot observe the state $x_{j_{k}}(t)$ ($k=1,\dots,p$) directly rather than giving a fixed communication topology, we can modify problem (15) as
$\min_{W,P}{}{}$ $\langle R,W\rangle+\gamma g(P)$ (17)
${\rm s.t.}{}{}$ $\mathcal{G}(W)\in\mathcal{K},$
$V_{1}WV_{2}^{\top}-P=0,$
$[V_{1}WV_{2}^{\top}]_{i_{k}j_{k}}=0,~{}k=1,\dots,p$
with $(i_{k},j_{k})\in[m]\times[n]$ for $k=1,\dots,p$. To the best of the authors’ knowledge, such special formulation, i.e., the design of a sparse feedback gain $K$ with forcing $[K]_{i_{k}j_{k}}=0$ for $k=1,\dots,p$, has not been reported in existing literature. It is worth to underline that problem (17) can be solved analogously by the proposed algorithms in Section 3, 4.
@@ 3 Two-Timescale Optimization Algorithm Framework
Let $\widetilde{W}={\rm vec}(W),\widetilde{P}={\rm vec}(P)$, then by the definition of indicator function $\delta_{\Gamma_{+}^{n}}(\cdot)$, (15) is equivalent to the following optimization problem
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\gamma g(\widetilde{P})$ (18)
${\rm s.t.}{}{}$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,~{}~{}i=1,\dots,M,$
@@
where $\widetilde{\Psi}_{i}={\rm vec}(\Psi_{i})$ and $\Gamma_{+}^{p}=\left\{{\rm vec}(W)\colon W\in\mathbb{S}_{+}^{p}\right\}$. Denoting $\mathcal{A},\mathcal{B},\widehat{\mathcal{A}}$ with
$\mathcal{A}=\begin{pmatrix}V_{12}^{\top}\otimes V_{11}\\ \vdots\\ V_{M2}^{\top}\otimes V_{M1}\\ V_{2}\otimes V_{1}\end{pmatrix}\doteq\begin{pmatrix}\widehat{\mathcal{A}}\\ V_{2}\otimes V_{1}\end{pmatrix},~{}~{}\mathcal{B}=\begin{pmatrix}0\\ \vdots\\ 0\\ -I_{mn}\end{pmatrix},$
optimization problem (18) is summarized in the following form
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $f(\widetilde{W})+h(\widetilde{P})$ (19)
${\rm s.t.}{}{}$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0,$
@@
$f(\widetilde{W})$ $=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}),$
$h(\widetilde{P})$ $=\gamma g(\widetilde{P}).$
To solve optimization problem (19), we briefly review the basic setup of primal-dual splitting method; see [39] for a comprehensive review. The following separable convex optimization problem is considered in [39]:
$\min_{x\in\mathcal{X},y\in\mathcal{Y}}F(x,y)=\bar{f}(x)+\bar{g}(y)~{}~{}{\rm s.t.}~{}Ax+By=b,$ (20)
where $\mathcal{X}\subseteq\mathbb{R}^{m}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n}$ are two closed convex sets, and $\bar{f}$ and $\bar{g}$ are two proper closed convex functions. Assume that $\bar{g}$ is closed and $\mu_{g}$-strongly convex with $\mu_{g}>0$ and $\bar{f}=f_{1}+f_{2}$ with closed convex $f_{2}$, $\mu_{f}$-strongly convex and $L_{f}$-smooth $f_{1}$ and $0\leq\mu_{f}\leq L_{f}<\infty$. Then, the following scheme is proposed to solve problem (20)
$u_{k}=(x_{k}+\alpha_{k}v_{k})/(1+\alpha_{k}),$ (21)
$d_{k}=\nabla f_{1}(u_{k})+A^{\top}\lambda_{k},$
$v_{k+1}=\underset{v\in\mathcal{X}}{\operatorname*{argmin}}\left\{f_{2}(v)+\langle d_{k},v\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|Av+Bw_{k}-b\right\|^{2}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|v-\widetilde{v}_{k}\right\|^{2}\right\},$
$x_{k+1}=(x_{k}+\alpha_{k}v_{k+1})/(1+\alpha_{k}),$
$\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(Av_{k+1}+Bw_{k}-b),$
$y_{k+1}=\mathbf{prox}_{\tau_{k}\bar{g}}^{\mathcal{Y}}(\widetilde{y}_{k}-\tau_{k}B^{\top}\bar{\lambda}_{k+1}),\quad\tau_{k}=\alpha_{k}^{2}/\eta_{g,k},$
$w_{k+1}=y_{k+1}+(y_{k+1}-y_{k})/\alpha_{k},$
$\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(Av_{k+1}+Bw_{k+1}-b),$
@@
$\widetilde{y}_{k}=y_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-y_{k}),$ (23)
$\widetilde{\eta}_{f,k}=\gamma_{k}+\mu_{f}\alpha_{k},$ (24)
$\widetilde{v}_{k}=\frac{1}{\widetilde{\eta}_{f,k}}(\gamma_{k}v_{k}+\mu_{f}\alpha_{k}u_{k}),$ (25)
and $\{(u_{k},d_{k},v_{k},x_{k},\bar{\lambda}_{k},y_{k},w_{k},\lambda_{k})\}$ is the iteration sequence. To emphasis, $\alpha_{k}>0$ is the step size and the parameter system is given by
$\frac{\theta_{k+1}-\theta_{k}}{\alpha_{k}}=-\theta_{k+1},$ (26)
$\frac{\gamma_{k+1}-\gamma_{k}}{\alpha_{k}}=\mu_{f}-\gamma_{k+1},$
$\frac{\beta_{k+1}-\beta_{k}}{\alpha_{k}}=\mu_{g}-\beta_{k+1}$
@@
is the continuous-time version of algorithm (21), where
$\mathcal{L}(x,y,\lambda)=\bar{f}(x)+\bar{g}(y)+\left\langle\lambda,Ax+By-b\right\rangle.$ (28)
It can be observed that differential inclusion (27) encompasses two Nesterov-type differential inclusions pertaining to primal variables $x$ and $y$, as well as a gradient flow about the dual variable $\lambda$. Besides, an alternative presentation of (27) reads as
@@
Therefore, (21) represents the semi-implicit Euler discretization of differential inclusion (29), which justifies its classification as a differential equation solver approach. We utilize this particular primal-dual splitting method instead of classic ADMM mainly because we can establish the relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function; this will be discussed in the following sections.
@@
We first consider the iteration of variable $\widetilde{P}\in\mathbb{R}^{mn\times 1}$. By (21), we have
$\widetilde{P}_{k+1}=\mathbf{prox}_{\tau_{k}h}^{\mathbb{R}^{mn\times 1}}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right)$
Denote
$\Omega^{k}={\operatorname*{vec}}^{-1}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right),$ (30)
$\rho_{k}=\frac{1}{\tau_{k}}$, $a_{k}=\frac{\gamma}{\rho_{k}}w_{ij}$, and let
$P_{ij;k+1}^{*}$ $\doteq\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\gamma w_{ij}|P_{ij}|+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right\}$ (31)
[…]
By Example 6.19 of [49], we can deduce that $\widetilde{P}_{k+1}=(P_{11;k+1}^{*},\dots,P_{mn;k+1}^{*})^{\top}$ which exhibits naturally sparsity structure.
@@ 4 Solving Inner Layer Optimization Problem
Algorithm (21) consists of two subproblems: updating $v_{k+1}$ and $y_{k+1}$. The solution to the latter can be obtained explicitly using equation (31), while the solution to the former will be discussed in this section.
By the update rule of $v_{k+1}$ in (21) and the definitions of $f_{2}$ and $b$, when employing the symbol system of this paper, the subproblem in (21) can be transformed to
$\min_{{\rm vec}(W)}{}$ $\langle d_{k},{\rm vec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}{\rm vec}(W)+\mathcal{B}w_{k}\right\|^{2}$ (32)
$+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|{\rm vec}(W)-\widetilde{v}_{k}\right\|^{2}$
${\rm s.t.}{}$ $\mathrm{vec}(W)\in\Gamma_{+}^{p},$
$\mathrm{vec}(\Psi_{i})\in\Gamma_{+}^{n},~{}i=1,\dots,M.$
By Assumption 3, for all $k$, problem (32) satisfies Slater’s condition and strong duality holds. Since $W,\Psi_{i}$ are symmetric matrices, there exist matrices $T_{1},T_{2}$ such that
${\rm svec}(W)$ $=T_{1}{\rm vec}(W)~{}{\rm with}~{}T_{1}T_{1}^{\top}=I_{\frac{p(p+1)}{2}},$
${\rm svec}(\Psi_{i})$ $=T_{2}{\rm vec}(\Psi_{i})~{}{\rm with}~{}T_{2}T_{2}^{\top}=I_{\frac{n(n+1)}{2}},$
where $\mathrm{svec}(\cdot)$ is the vectorization operator on $\mathbb{S}^{n}$ defined by
@@
Inversely, there exist matrix $D_{1},D_{2}$ such that
${\rm vec}(W)=D_{1}{\rm svec}(W),~{}~{}{\rm vec}(\Psi_{i})=D_{2}{\rm svec}(\Psi_{i}).$
@@
We partition the identity matrix $I_{N}$ ($N=n(n-1)/2$) as follows
$I_{N}=(u_{11},u_{21},\dots,u_{n1},u_{22},\dots,u_{n2},\dots,u_{nn}).$
@@
with $E_{ij}=e_{i}e_{j}^{\top}$. By Lemma 3.2 of [50], $L^{\top}=L^{\dagger}$, and naturally $LL^{\top}=I_{N}$.
Inversely, there exist matrix $D_{1}\in\mathbb{R}^{p^{2}\times\frac{p(p+1)}{2}}$ and $D_{2}\in\mathbb{R}^{n^{2}\times N}$ such that
${\rm vec}(W)=D_{1}{\rm svec}(W),~{}~{}{\rm vec}(\Psi_{i})=D_{2}{\rm svec}(\Psi_{i}).$
@@
Then, by utilizing $\mathrm{svec}(\cdot)$ operator and denoting ($\widetilde{\Gamma}_{+}^{n}$ is defined analogously)
@@
$\min_{{\rm svec}(W)}{}$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}$ (33)
$+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$
${\rm s.t.}{}$ ${\rm svec}(W)\in\widetilde{\Gamma}_{+}^{p}$
${\rm svec}(\Psi_{i})\in\widetilde{\Gamma}_{+}^{n},~{}i=1,\dots,M,$
where ${\rm svec}(\Psi_{i})=T_{2}{\rm vec}(\Psi_{i}).$ The primal problem of (33) presents a level of difficulty that warrants exploration of its dual problem. The Lagrangian of (33) is defined as
$\mathcal{L}({\rm svec}(W);X)$ (34)
$=$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}$
$+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}-\left\langle X_{0},{\rm svec}(W)\right\rangle$
$+\left\langle(X_{1}+\dots+X_{M}),T_{2}(V\otimes V){\rm vec}(Q)\right\rangle$
@@
$\tilde{d}_{k}=D_{1}^{\top}d_{k}-X_{0}$
$~{}~{}~{}~{}~{}~{}+D_{1}^{\top}\sum_{i=1}^{M}\left[(V\otimes VF_{i})^{\top}+(VF_{i}\otimes V)^{\top}\right]T_{2}^{\top}X_{i}$
$~{}~{}~{}\doteq\mathcal{L}_{k}(X_{0},X_{1},\dots,X_{M}),$
$\sigma_{1}=\frac{\alpha_{k}}{2\theta_{k}},~{}\sigma_{2}=\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}},~{}\tilde{b}_{k}=\mathcal{B}w_{k},$